If p^n=cos^n x+sin^n x, then 2p^6-3p^4+1 is equal to what

Grade:10

1 Answers

Arun
5 years ago
Dear student
 
p^6 = cos ^6 x + sin^6 x = (cos^2 x + sin^2 x) (cos^4 x + sin^4 x – cos^2 x sin^2 x) = 1 – 3 s^2 c^2
 
p^4 = 1 – 2 s^2 c^2
 
hence 
2 p^6 – 3 p^4 + 1 = 0